OpenStudy (anonymous):

Molly went to the beach and collected a bag of seashells. She kept half of her collection. She gave the other half to Beth and Jack. Jack received three times as many seashells as Beth. If Beth received 8 seashells, how many seashells did Molly collect at the beach? 64 56 32 24

OpenStudy (ganpat):

Molly got a bag containing x sea shells.. 0.5x = B + J J = 3B B = 8, so jack = 3 (8 ) = 24.. Jack + Berth = 24 + 8 = 32.. Molly gave half to J & B .. so Molly collected = 2 (32) = 64... sea shells...

OpenStudy (anonymous):

thank you :D

OpenStudy (ganpat):

W/c :)